OpenStudy (anonymous):

Solve the initial value problem: x(dy/dx)+y(x) = 9y(x)^(2), y(1) = -1

OpenStudy (anonymous):

\[x \frac{ dy }{dx } + y(x) = 9y(x)^2 , y(1) =-1\] Just the equation a little neater

OpenStudy (anonymous):

I would solve for dy/dx and intergrate.

OpenStudy (unklerhaukus):

the ex's int he brackets are just indicating the independent variable right?\[x \frac{ \text dy }{\text dx } + y = 9y^2 , \]

OpenStudy (anonymous):

and we know y(1) = -1.

OpenStudy (anonymous):

I believe so, that's just how its been written. I'll try solving for dy/dx and integrating then,

OpenStudy (anonymous):

So I would sub those in.

hartnn (hartnn):

u realize that you can separate the variables easily here ?

OpenStudy (anonymous):

\[\frac{ dy }{ dx } = \frac{ 9y(x)^2-y }{ x }\] Not too familiar with these problems, but basically i need to integrate that, no?

hartnn (hartnn):

u bring all terms of one variable on one side of = sign, like this : \(\large \frac{1}{9y^2-y}dy=\frac{1}{x}dx\) then integrate

hartnn (hartnn):

can u integrate both sides now ?

OpenStudy (anonymous):

doing that now

hartnn (hartnn):

take your time :)

OpenStudy (anonymous):

think i got it, left an x over the y side so i confused my self. \[1 = \ln(1-9y)-\ln(y)\]

hartnn (hartnn):

u integrated y-variable correctly , but what about \(\int (1/x)dx\) its not =1

hartnn (hartnn):

i meant u should get something like this : \(\ln x=ln(1-9y)-lny+c\) then use logarithmic properties to simplify

OpenStudy (anonymous):

Okay, catching on think i figured out what i did wrong when i integrated last anyway.

OpenStudy (anonymous):

Maybe y = \[\frac{ 1 }{ x+9 }\] May be a final solution, just simplifying the equation?

hartnn (hartnn):

ln cx= ln |(1-9y)/y| cxy= (1-9y) now use y(1) = -1 to find c.

hartnn (hartnn):

u got this simplification ? ---->ln cx= ln |(1-9y)/y|

OpenStudy (anonymous):

yup, \[C(-1)(1)=1-9(-1)\] \[-C=10\] \[C=-10\] \[10xy=(1-9y)\]

hartnn (hartnn):

-10xy =1-9y or 10xy-9y+1=0

OpenStudy (anonymous):

ahh, forgot the negative, think i have enough to try a few more of these questions any way. Thank you. :)

hartnn (hartnn):

welcome ^_^